# Dense Backpropagation Improves Training for Sparse Mixture-of-Experts

### v1

@@ 1 Introduction
Figure 1: Pretraining Benchmarks for Default MoE Our Default MoE improves MoE pretraining by estimating the dense gradient of the router. After training for 320B tokens on FineWeb-Edu (Lozhkov et al., 2024), our 2 billion parameter Default MoE outperforms a standard TopK MoE on various pretraining benchmarks. All evals are run with the lm-eval-harness (Gao et al., 2024).
@@ 4 Evaluation
We describe the experimental setup in Section 4.1, show that our Default MoE method speeds up training by $15\%$ compared to the TopK baseline in Section 4.2, and then show how our method compares to the baseline as we ablate virtually every parameter in the experimental setup in Section 4.3. We finish by concluding that our method does not significantly impact throughput or the memory footprint of training, so the improvements we observe are -as far as we can tell- a free lunch.
@@ 4.1 Experimental Setup
We train a standard MoE with 8 experts. The model has a hidden dimension of $1024$, with 2B total parameters. When doing $K=1$ top-K routing, only 1 expert at each layer is activated, and about $500M$ of the parameters are active. We ablate the model architecture, hidden dim, number of total experts, and number of active experts in Section 4.3. We train on FineWeb (Penedo et al., 2024) with the Llama3 tokenizer (Llama 3 Team, 2024). We use SwiGLU (Shazeer, 2020) MLPs following Llama (Touvron et al., 2023), $16$ attention heads with dimension of $64$, LayerNorm (Ba et al., 2016) and RoPE (Su et al., 2023).
Hyperparameters. We use the initialization from  Wang et al. (2022) for residual branch merge layers and the initialization from Nguyen & Salazar (2019) for all other layers. We use the AdamW optimizer (Loshchilov & Hutter, 2019). We set the maximum learning rate to $7\times 10^{-4}$ following a hyperparameter sweep, the minimum learning rate to $7\times 10^{-5}$, and use a standard cosine decay schedule. We use a sequence length of $2048$ and a global batch size of $1024$, resulting in a global token batch size of $2^{21}$.
@@
Implementation. We train with the gpt-neox library (Andonian et al., 2023) integrated with Megablocks (Gale et al., 2022) and augmented with a number of Triton kernels that speed up training. We have open-sourced our implementation already for reproducibility, and will include a link here upon publication.
@@ 4.2 Main Results
Pretraining Benchmarks. Our main result compares our Default MoE, which performs a dense update of the router weights by approximating the dense gradient, to a baseline Top-K MoE, both with $K=1$ active experts. After training both our Default MoE and a TopK MoE for 320 billion tokens, we evaluate both models on a variety of common pretraining benchmarks in Table 1. We plot benchmarks for both models in Figure 1. Our Default MoE outperforms TopK MoE across the board, with an average benchmark improvement of 2.8%.
@@ 4.3 Ablations
We now ablate every setting in our experimental setup, from the model architecture to the learning rate.
@@
Figure 5: Ablating the hidden dimension. We perform an ablation study across varying hidden dimensions for both the Default MoE and a Top-K baseline approach. We plot the perplexity across 3 hidden dimensions: 512, 1024 and 2048. Notably, the Default MoE consistently outperforms the Top-K baseline across all model sizes, demonstrating lower perplexity at every scale. We compare the training efficiency of Default MoE and Top-K by measuring the number of tokens required to reach a target perplexity. Since Default MoE uses an Exponential Moving Average (EMA) of previous expert outputs to supplement the router during training, it requires fewer tokens to reach the same perplexity as Top-K.
Default MoE Remains Superior As We Increase the Model Size.  Figure 5 compares the two methods as we increase the size of the model from $557M$ total parameters at a hidden dimension of $512$, to $7.33B$ total parameters when the hidden dimension is $2048$. Our method outperforms the baseline across all model sizes. All other results in the paper use a hidden dimension of $1024$ because this model is fairly easy to train with just data parallelism.
@@ 4.4 Efficiency
Our Method Does Not Significantly Reduce Throughput. In Table 2 we report the throughput in samples per second and of TopK and our method while training the standard 8-expert MoE on a single GPU node. Our method is quite lightweight, because we just need to update an EMA in the forward pass and then update the router gradients in the backward pass. However, for the 2B model we train, this is still nearly a $2\%$ overhead; less than the $9\%$ speedup we observe over the TopK baseline in Figure 4, but still non-zero. We find that the overhead of our method decreases as we increase the hidden size. The $2B$ model we train has $1024$ hidden size, $2048$ hidden size is an $8B$ model, etc. As the hidden size increases, the proportion of time spent in the MLP matmuls increases, and the overhead of our model is no longer significant compared to the total MoE layer runtime. For even just an 8B model, our overhead is near-zero; and of course, this is almost 100×100$\times$100 × smaller than production MoEs (DeepSeek-AI Team, 2024b). We are confident that our lightweight EMA method does not reduce throughput at scale.

### v2

@@ Abstract
Figure 1: Pretraining Benchmarks for Default MoE. Our Default MoE improves MoE pretraining by providing a default vector for unactivated expert activations. After training for 320B tokens on FineWeb-Edu (Lozhkov et al., 2024), our 2B parameter Default MoE outperforms a standard TopK MoE on various pretraining benchmarks. All evals are run with the lm-eval-harness (Gao et al., 2024).
@@ 4 Evaluation
We describe the experimental setup in Section 4.1, show that our Default MoE outperforms the TopK baseline in Section 4.2, and then show how our method compares to the baseline as we ablate virtually every parameter in the experimental setup in Section 4.3. We finish by concluding that our method does not significantly impact throughput or the memory footprint of training, so the improvements we observe are -as far as we can tell- a free lunch.
@@ 4.1 Experimental Setup
Model Architectures. We use NcK to refer to a model with $N$ total experts and $K$ active experts. For our main results in Figure 1Figure 4Figure 4 we train a standard MoE with 8 experts. All of our models have 1.96 billion total parameters, 366 million of which are non-MoE parameters. This leaves 1.6B MoE parameters, and the number of active parameters depends on sparsity. For example, a 32 expert MoE with 2 active experts has sparsity 1/16, so 100M of the MoE parameters are active i.e. the 32c2 MoE has 466M active parameters. All parameter counts are specified in Table 1.
MoE Config. 8c1 8c2 32c1 32c2 32c4
Active Params. 565M 764M 416M 466M 565M
Table 1: Active parameter counts for each MoE configuration; the total parameter count is 1.96B.
We ablate the model architecture, hidden dim, number of total experts, and number of active experts in Section 4.3.
Dataset. We train on FineWeb-Edu (Lozhkov et al., 2024) (for Figure 1 and Figure 4) and FineWeb (Penedo et al., 2024) (for all other plots and results) with the Llama3 tokenizer (Llama 3 Team, 2024). We use SwiGLU (Shazeer, 2020) MLPs following Llama (Touvron et al., 2023), $16$ attention heads with dimension of $64$, LayerNorm (Ba et al., 2016) and RoPE (Su et al., 2023).
Hyperparameters. We use the initialization from  Wang et al. (2022) for residual branch merge layers and the initialization from Nguyen & Salazar (2019) for all other layers. We use the AdamW optimizer (Loshchilov & Hutter, 2019). All hyperparameters for the model that produces Figure 1 can be found in our config file, and all hyperparameters are the same between the baseline and our DefaultMoE.
@@
Implementation. We train with the gpt-neox library (Andonian et al., 2023) integrated with Megablocks (Gale et al., 2022) and augmented with Triton kernels from (Hsu et al., 2024). We open-source our training code.
@@ 4.2 Main Results
Pretraining Benchmarks. Our main result compares our Default MoE, which performs a dense update of the router weights by approximating the dense gradient, to a baseline Top-K MoE, both with $K=1$ active experts. After training both our Default MoE and a TopK MoE for 320 billion tokens on FineWeb-Edu (Lozhkov et al., 2024), we evaluate both models on a variety of common pretraining benchmarks in Table 2. We plot benchmarks for both models in Figure 1. Our Default MoE outperforms TopK MoE across the board, with an average benchmark improvement of 2.8%.
@@
Figure 4: Default MoE beats a tuned TopK baseline on FineWeb. Our method reaches the target perplexity of 18 about 15% faster than the baseline TopK method, without introducing any additional overhead.
Pretraining on FineWeb.
In Figure 1 and Figure 4 we pretrain on FineWeb-Edu (Lozhkov et al., 2024). We also pretrain models on FineWeb (Penedo et al., 2024), which is less curated. The rest of the plots in this paper (our ablations) are on FineWeb, in order to enforce that on some level we are not simply overfitting to the pretraining dataset that we use to collect benchmark scores. In Figure 4 we show that when using the learning rate of $7\times 10^{-4}$ that we sweep in Figure 8, our DefaultMoE outperforms the baseline TopK. Therefore, our improvements are not due to any unfair hyperparameter tuning when compared to the baseline.
@@ 4.3 Ablations
We now ablate every setting in our experimental setup, from the MoE config. to the learning rate.
@@
Figure 6: Ablating the hidden dimension. We perform an ablation study across varying hidden dimensions for both the Default MoE and a Top-K baseline approach. We plot the perplexity across 3 hidden dimensions: 512, 1024 and 2048. Notably, the Default MoE consistently outperforms the Top-K baseline across all model sizes, demonstrating lower perplexity at every scale.
Default MoE Remains Superior As We Increase the Model Size.  Figure 6 compares the two methods as we increase the size of the model from $557M$ total parameters at a hidden dimension of $512$, to $7.33B$ total parameters when the hidden dimension is $2048$. Our method outperforms the baseline across all model sizes. All other results in the paper use a hidden dimension of $1024$.
@@ 4.4 Efficiency
Our Method Does Not Significantly Reduce Throughput. In Table 3 we report the throughput in samples per second and of TopK and our method while training the standard 8-expert MoE on a single GPU node. Our method is quite lightweight, because we just need to update an EMA in the forward pass and then update the router gradients in the backward pass. However, for the 2B model we train, this is still nearly a $2\%$ overhead; less than the $9\%$ speedup we observe over the TopK baseline in Figure 4, but still non-zero. We find that the overhead of our method decreases as we increase the hidden size. The 2B model we train has $1024$ hidden size, $2048$ hidden size is an 2B model, etc. As the hidden size increases, the proportion of time spent in the MLP matmuls increases, and the overhead of our model is no longer significant compared to the total MoE layer runtime. For even just an 8B model, our overhead is near-zero; and of course, this is almost 100×100$\times$100 × smaller than production MoEs (DeepSeek-AI Team, 2024b). We are confident that our lightweight EMA method does not reduce throughput at scale.